# Network-aided Efficient LLM Services With Denoising-inspired Prompt Compression

### v3

@@ 5 Numerical Results
Fig. 7 plots the performance when running with the LongBench dataset and compares the scores gained by no compression method, a single compression method with a single 16x compression ratio, and an iterative compression method with an iterative 4 times 2x compression ratio. The comparison is conducted from the perspective of both the average QA F1 score [37] and the Rouge-L score [38] using the Multi-news and Govreport data types. The F1 score evaluates question-answering systems, with higher values indicating better performance. A score of 1 indicates perfect classification, while a score of 00 means complete misclassification. It balances precision and recall. ROUGE-L assesses text generation quality by comparing outputs to reference texts. The ’L’ stands for Longest Common Subsequence (LCS), which captures matching sequences in order, though not necessarily consecutively. It is observed from the figure that for the multi-news category, the no compression, single compression, and iterative compression methods achieve similar scores. In the govreport category, the iterative compression method achieves higher scores in both the F1 score and ROUGE-L score.
Figure 8: The illustration of network throughput improvement and service time reduction. Network throughput improvement corresponds to the reduction in transmission time, and service time includes the compression time, transmission time, and inference time as shown in Fig. 1.
Fig. 8 plots the network throughput improvement and service time reduction with $5$ data examples randomly sampled from the MeetingBank-transcript dataset. The network throughput improvement is the reduction in transmission time of the compressed prompt, showing the decrease in transmission time using the proposed denoising-inspired prompt compression method compared to the single compression method. The service time is the sum of the compression time, transmission time and inference time as shown in Fig. 1. The ratio in the figure shows the reduction of service time using the proposed denoising-inspired prompt compression method compared with the single compression method. It shows that the proposed denoising-inspired prompt compression method can consistently enhance network throughput while simultaneously reducing service time. The results indicate that this approach effectively optimizes data transmission efficiency, ensuring faster response times without compromising performance. This consistent improvement highlights the method’s robustness in handling different data types and network scenarios, making it a reliable solution for efficient network communication.
@@ 6 Conclusion and Future Direction
We proposed a novel joint power and prompt optimization framework for wireless network-aided accelerated LLM services. Additionally, inspired by denoising diffusion models, we designed a denoising-inspired prompt compression scheme to further enhance system performance. The proposed approach delivers high service quality with a relatively low BEP while keeping power consumption within limits. Numerical results show that the proposed algorithm demonstrates stable convergence, highlighting its suitability for practical deployment in LLM service systems.
Several promising directions for future research emerge from this work. While our current implementation treats the denoising-inspired prompt compression scheduling type and the iteration count as fixed parameters to demonstrate performance improvements, these could be incorporated as optimization variables in future studies. Different text types may benefit from different compression schedules, and the number of iterations, being directly linked to computing resources and system performance, could be integrated into the optimization problem. The ability to optimize both power and prompt compression makes this approach well-suited for resource-constrained environments, such as autonomous systems, industrial IoT, and smart home networks. These domains require efficient data transmission and minimal latency, aligning well with our proposed method. Additionally, integrating adaptive prompt compression with real-time processing in edge computing can enhance system performance while maintaining power efficiency. Furthermore, the deployment of SLMs at edge devices, facilitated by their small parameter count, opens up new research opportunities regarding their role in communication networks and their interaction with LLMs.

### v4

@@ 5.2 Experiments Performance Analysis
Fig. 7 compares the task performance of three compression strategies, i.e., no compression, JPPO with single-round 16×16$\times$16 × compression, and JPPO++ with iterative 4×2×4$\times$ 2$\times$4 × 2 × compression, on the LongBench dataset. Results are reported on both the Multi-news and GovReport categories using the average Question-Answering (QA) F1 score and ROUGE-$L$ score. The no-compression strategy achieves the highest performance across all samples. In the Multi-news category, as shown in Fig. 7 (a), where contextual coherence is less critical, the gap between compressed and uncompressed prompts is relatively small. However, in the GovReport category, as shown in Fig. 7 (b), which contains longer and more information-dense documents, the difference becomes more pronounced. Despite both JPPO and JPPO++ operating under the same 16×16$\times$16 × compression constraint, JPPO++ consistently yields higher F1 and ROUGE-$L$ scores, demonstrating better retention of key information through iterative refinement. These results highlight the practical advantage of JPPO++. It enables aggressive compression with minimal degradation of task performance. Even though uncompressed input naturally yields the best performance, JPPO++ offers a viable trade-off for resource-constrained deployments, maintaining task utility while significantly reducing prompt length.
@@ 6 Conclusion and Future Direction
We proposed JPPO++, a joint optimization framework that integrates denoising-inspired prompt compression with wireless transmission power control to support mobile LLM services. By leveraging a lightweight SLM for iterative prompt compression and a DRL-based policy for adapting compression ratio and power allocation, JPPO++ enabled aggressive compression while preserving response fidelity under resource constraints. The framework demonstrated stable convergence during training, and experimental results showed that it consistently reduced both service time and transmission overhead compared to single-compression baselines. Specifically, JPPO++ achieved up to $46.5\%$ service time reduction and maintained competitive performance on downstream QA and summarization tasks, even under a 16×16$\times$16 × compression setting. These findings demonstrate the practical value of JPPO++ in enabling scalable and low-latency mobile LLM services.
There are several directions for further research:
  • Optimizing Compression Schedules : The current JPPO++ uses a fixed denoising schedule and iteration count. Future work may treat both as decision variables. Different input types may benefit from tailored schedules, and the number of iterations directly affects computational cost and compression quality. Including these factors in the optimization can further improve performance across diverse tasks.
  • Edge Deployment and Distributed Control: SLMs can be deployed at edge devices to perform real-time compression before wireless transmission. Future studies may explore their interaction with LLMs in hierarchical settings. Additionally, transitioning from centralized to distributed optimization, where multiple edge nodes collaboratively manage compression and power allocation, can enhance system scalability and resilience.
  • Personalized and Adaptive Service Provisioning: Incorporating user-specific demands and heterogeneous network conditions into the joint optimization can enable context-aware service delivery. This is particularly relevant in autonomous systems and smart home networks, where latency and efficiency requirements vary significantly.